OpenStudy (anonymous):

A ship is traveling around the outside rim of an island surveying a good place to land. If it is 10 miles from a lighthouse at the center of the island and 45° from due north, how far has the ship traveled from the most northern tip of the island to the nearest tenth? @ganeshie8

OpenStudy (anonymous):

@cwrw238

OpenStudy (secret66):

Circumference of island = 2*pi*10 miles. That embraces 360 degrees. 45 degrees change in bearing = (45/360)*2*pi*10 miles = (5/2)*pi miles

OpenStudy (anonymous):

78?

OpenStudy (anonymous):

@secret66

OpenStudy (secret66):

What's (5/2)*3.14?

OpenStudy (anonymous):

7.85

OpenStudy (secret66):

Yeah...